# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### v3

@@ 5.3 Ablation Studies
Effect of the Recursive Redirection Strategy. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab. 4 where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase. One naive baseline is to remove the recursive seed sampling component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Therefore, the performance is limited. The proposed method achieves the best performance with less than 10% running time increase on the BvEM-Macaque dataset. Unfortunately, it significantly increases the computation cost. Compared to our method, DenseSAM’s running time is 5.97 times longer since it needs to segment 3 planes for every tracking step. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
SAM [15] 77.65 66.03 55.48 1535s 615M
MobileSAM [35] 78.41 74.97 62.14 335s 9.66M
Quality Analysis. We delve deeper into the plane dynamics in Fig. 11 and 11. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Tab.3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis in the test volume. In Fig. 11, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Other Ablations. To further improve the efficiency, we use MobileSAM instead of the original SAM in practice. As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig. 8 illustrates the performance with varying hyperparameters $\eta$ and $\tau$.
Figure 12: The failure cases of the proposed method. Three prediction and ground truth pairs are shown.
@@ 5.4 Failure Cases
Despite our method’s significant progress, it still encounters challenges in some scenarios, as illustrated in Figure 12. The first issue is the seed problem, where the presence of a false positive seed can lead to erroneous predictions. Secondly, SAM could also fail to segment the mask. Lastly, tracking failure tends to occur at blood vessel conjunction points. These instances highlight the complexity of tracking in such intricate vascular networks and suggest the need for further refinements in our approach to address these specific failure cases effectively.

### v4

@@ 5.3 Ablation Studies
Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
SAM [15] 77.65 66.03 55.48 1535s 615M
MobileSAM [36] 78.41 74.97 62.14 335s 9.66M
SAM-based tracking module. We examine the effect of different SAM variants, i.e., MobileSAM, on the TriSAM performance. As shown in Tab. 4, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.
Turning point sampling module. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab. 5 where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase. One naive baseline is to remove the recursive seed sampling component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Therefore, the performance is limited. The proposed method achieves the best performance with less than 10% running time increase on the BvEM-Macaque dataset. Unfortunately, it significantly increases the computation cost. Compared to our method, DenseSAM’s running time is 5.97 times longer since it needs to segment 3 planes for every tracking step. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
@@
Tri-plane dynamics. We delve deeper into the plane dynamics in Fig. 10 and 10. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Tab.3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis in the test volume. In Fig. 10, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Figure 11: Ablation study results on hyperparameters $\eta$ and $\tau$.
Hyperparameters. We compute the accuracy of TriSAM with varying hyperparameters for seed generation ($\eta$) and SAM-based tracking module ($\tau$). As shown in Fig. 11, TriSAM’s performance is robust within the reasonable range of hyperparameters.
@@ 5.4 Failure Cases
The main source of error in the proposed TriSAM framework is the SAM-based tracking module, where SAM can not produce accurate 2D mask given the prompt. For the false negative case (Fig. 12a), SAM predicts an empty mask when the shape and appearance of the blood vessels are complex (e.g., conjunction point) for all three axes. For the false positive case (Fig. 12b), SAM falsely segment the neighboring region with the blood vessel due to the similar image appearance.
@@ 6 Conclusion
Figure 12: Failure cases of the SAM-based tracking module.